Find the course and distance from a position 37⁰ 01’ N 9⁰ 00’W to a position 36⁰ 11’N
6⁰ 02’W.
1
Expert's answer
2020-10-26T19:12:25-0400

"φ_A=37^{\circ}01'" N; -the geographical latitude of position A

"λ_A=9^{\circ}00'" W; -the geographical longitude of position A

"φ_B=36^{\circ}11'" N; -the geographical latitude of position B

"λ_B=6^{\circ}02'" W; -the geographical longitude of position B

"\Delta σ=arccos(sin(φ_A)sin(φ_B)+"

"+cos(\u03c6_A)cos(\u03c6_B)cos(\u03bb_B-\u03bb_A))="

"=arccos(sin(37^{\circ}01')sin(36^{\circ}11')+"

"+cos(37^{\circ}01')cos(36^{\circ}11')cos(6^{\circ}02'-9^{\circ}00'))="

"=arccos(0.6020\cdot 0.5904+"

"+0.7985\cdot 0.8071\cdot cos(-2^{\circ}58'))="

"=arccos(0.3554+0.6445\cdot 0.9986)="

"=arccos(0.9990)=2.52^{\circ}" - the central angle between A and B;

the great-circle distance between two points "AB=60\\cdot2.52^{\\circ}=151.2\u202c" nautical miles,

or "AB=R\cdot2.52^{\circ}\frac{π}{180^{\circ}}=6371\cdot0.0440=280.3" km, where R=6371 km - the mean earth radius.


"α=arccos(\frac{sin(φ_B)-sin(φ_A)\cdot cos(\Delta σ)}{cos(φ_A)\cdot sin(\Delta σ)})="

"=arccos(\frac{sin(36^{\circ}11')-sin(37^{\circ}01')\cdot cos(2.52^{\circ})}{cos(37^{\circ}01')\cdot sin(2.52^{\circ})})="

"=arccos(\frac{ 0.5904-0.6020\cdot 0.9990}{0.7985\cdot 0.0439})="

"arccos(-0.3137)=108.3^{\circ}" - bearing from point A to point B


Answer: the distance between positions is about 151.2 nautical miles (280.3 km),

course (bearing from point A to point B) is about 108.3⁰.